# A Cross-Font Image Retrieval Network for Recognizing Undeciphered Oracle Bone Inscriptions

### v1

@@ III-D Ablation Studies
As shown in Tables III and IV, we conduct several ablation studies to further demonstrate the effectiveness of the components of CFIRN.
Effectiveness of MFI. To investigate the contribution of the multiscale feature integration (MFI) module to feature extraction, we conducted corresponding ablation studies. The results indicate that the complete model equipped with the MFI outperforms the version without MFI across all evaluation metrics, especially in Recall@1 and AP. These findings underscore the significant role of the MFI in enhancing the model’s performance.
Effectiveness of MRC. We conducted studies to further validate the enhancing effect of the multiscale refinement classifier (MRC) on model performance. The experimental results demonstrated that the full model, equipped with the MRC module, outperformed the version without it noticeably. This is because the classifier is closely tied to the training loss and functions as more than just a feature refinement component.
Mixed-scale Input Scheme. We have observed that different combinations of input sizes have a partial impact on model performance. The experimental results indicate that the combination of using both the original and reduced sizes (0.5x, 1.0x) yielded the best performance across several evaluation metrics for image retrieval, suggesting that well-chosen scaling factors can enhance the model’s robustness and precision.

### v2

@@ III-D Ablation Studies
As shown in Tables IV and V, we conducted comprehensive ablation studies to verify the effectiveness of CFIRN components and configurations.
Effectiveness of MFI. To evaluate the contribution of the multiscale feature integration (MFI) module to feature extraction, we performed ablation studies as shown in rows 1 and 2 of Table IV. The results demonstrate that the full model incorporating MFI consistently outperforms the version without it across all metrics, particularly in Recall@1 (69.32% vs. 68.39%) and AP (55.89% vs. 54.94%). These results highlight the critical role of MFI in improving model performance.
Effectiveness of MRC. To isolate and evaluate the contribution of the multiscale refinement classifier (MRC), we compared the performance of models with and without the MRC module (rows 2 and 3 in Table IV). The results reveal significant performance improvements in Recall@5 (78.10% vs. 67.05%) and Recall@10 (81.51% vs. 71.32%) when MRC is included. This improvement arises because the MRC not only refines features but also directly influences the training loss, acting as a crucial component of the model.
Mixed-scale Input Scheme. We explored the influence of mixed-scale input combinations on model performance, as detailed in Table V. The results reveal that using both reduced (0.5x) and original (1.0x) input resolutions achieves the best performance, particularly in Recall@1 (69.32%) and AP (55.89%), compared to other strategies. This suggests that an appropriate mix of scaling factors can effectively enhance model robustness and retrieval precision.
Figure 6: Retrieval results and corresponding modern Chinese characters of the real-world undeciphered OBI characters.
III-E Real-World Undeciphered Character Analysis
For the decipherment of real-world undeciphered OBI characters, we selected the BI as the gallery font, given its closest temporal proximity and the completeness of its collected data. To maximize the utility of limited data resources, we utilized all OBI characters and BI characters that share annotated categories for training. In testing, undeciphered OBI characters retrieved all annotated BI characters (3354 classes) to produce the final decipherment results.
As shown in Fig. 6, we illustrate a series of challenging retrieval samples and their corresponding deciphered modern Chinese characters. The retrieval results visually demonstrate the robustness and precision of our model in handling the complex morphological variations inherent in undeciphered OBI characters. We hope these decipherment results can offer valuable insights for paleographic researchers.